An effective rate of 29,61% corresponds to a nominal rate, compounded weekly, of

1.     29,53%.

2.     29,61%.

3.     34,35%.

4.     26,00%.

5.     none of the above.


1
Expert's answer
2021-05-17T14:05:56-0400


Let effective interest rate=i,

nominal interest rate is j,

"i=100((\frac{1+j}{m})^m-1)",

29.61=100((1+r/52)52-1)).

Put r=26,

29.61=100((1+26/52)52-1)),

29.61=29.61.

LHS=RHS.

Thus, the option 4 is right.